Nickel Inc. bought $500,000 of 3-year, 9% bonds as an investment on December 31, 2017 for $545,000. The investment receives interest annually and Nickel uses straight-line amortization. On May 1, 2018, the issuer retired $100,000 of the bonds at 112. As a result of the retirement, Nickel will report a
Answer #1

Value of $100000 bonds = 545000*100000/500000 = $109000

Amortization of premium in 4 months = (109000-100000)*4/36 = $1000

Carrying value on May 1, 2018 = 109000 - 1000 = $108000

Retirement value = 100000*1.12 = $112000

Gain on retirement = 112000 - 108000 = $4000